OpenStudy (anonymous):

Simplify the expression. Write the answer using scientific notation. 8(8.8 x 1012) I put A. and still got it wrong. A. 70.4 x 1012 B. 70.4 x 1024 C. 70.4 x 1013 D. 1.68 x 1013

OpenStudy (zepp):

http://puu.sh/KRU1 Never forget to put the parentheses!

OpenStudy (zepp):

Wait, is it \(\large 8.8(8.8*10^{12})\)?

OpenStudy (anonymous):

yesssss

OpenStudy (zepp):

Strange, A) B) C) aren't even scientific notation :| @AccessDenied

OpenStudy (accessdenied):

Either there is missing information here, or there is no correct answer from the selection. That's how I see it.

OpenStudy (anonymous):

so what do i do? contact my school?

OpenStudy (zepp):

E) \(\large 7.74 * 10^{13}\)

OpenStudy (anonymous):

i just think that when you copy/pasted, it didn't make the copy the powers of 10 properly...

OpenStudy (anonymous):

@zepp are you sure????

OpenStudy (anonymous):

*it didn't copy...

OpenStudy (zepp):

Well, calculators don't lie :s

OpenStudy (anonymous):

it wasn't one of the anwsers

OpenStudy (anonymous):

hmm.. you're right... none of them are scientific notation....(except for the last one which is not correct)

OpenStudy (anonymous):

actually using my abacus, i got \(\large 7.04 \times 10^{13} \)

OpenStudy (anonymous):

@dpaInc im gonna try it

OpenStudy (anonymous):

wait... what is the correct problem: \[\huge 8.8(8.8\times 10^{12}) \] or \[\huge 8(8.8\times 10^{12}) \]

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

ok... my abacus does not lie either then....:)

OpenStudy (anonymous):

mk thank you!

OpenStudy (anonymous):

yw...:)